# No-prior Bayes reIMagined: probabilistic approximations of possibilistic inferential models

### v2

@@ 3 Inner probabilistic approximations
A well-known characterization (e.g., Couso et al., 2001; Destercke and Dubois, 2014) of the credal set for general possibility measures applied to the present case says
$\mathsf{Q}_{x}\in\mathscr{C}(\overline{\Pi}_{x})\iff\mathsf{Q}_{x}\{C_{\alpha}(x)\}\geq 1-\alpha\;\;\text{for all $\alpha\in[0,1]$},$ (7)
@@
where $C_{\alpha}(x)$ is as defined in (5) with $\pi_{x}$ the contour corresponding to $\overline{\Pi}_{x}$. Validity of the IM implies that $C_{\alpha}(x)$ is a $100(1-\alpha)$confidence set, so, since the elements $\mathsf{Q}_{x}$ in the credal set assign at least probability $1-\alpha$ to the $100(1-\alpha)$confidence seet $C_{\alpha}(x)$, there’s good reason to call these elements confidence distributions. This definition of confidence distributions agrees with that given in Taraldsen, (2021) and generalizes those commonly found in the statistics literature (Xie and Singh, 2013; Schweder and Hjort, 2013; Cunen et al., 2020) that—for reasons discussed in Section 4.3—focus mainly on scalar parameters and the corresponding univariate confidence distributions. This connection between confidence distributions and the credal set associated with the possibilistic IM was first made in Martin, 2023a , and I’ll revisit the specific results in Section 4.1.
@@
Refinements to the characterization in (7) can be given. For example, Wasserman, 1990b (, Theorem 2.1) characterizes the credal set associated with a belief function as a collection of suitable mixture distributions. The result below gives a similar mixture characterization but for general possibility measures, including for the IM output $\overline{\Pi}_{x}$.
Theorem 1 (Martin, 2025).
@@
Let $\pi_{x}$ be the contour associated with the possibility measure $\overline{\Pi}_{x}$, and set $C_{\alpha}(x)=\{\theta:\pi_{x}(\theta)\geq\alpha\}$ as in (5). Then $\mathsf{Q}_{x}\in\mathscr{C}(\overline{\Pi}_{x})$ if and only if
$\mathsf{Q}_{x}(\cdot)=\int_{0}^{1}\mathsf{K}_{x}^{\beta}(\cdot)\,\mathsf{M}_{x}(d\beta),$ (8)
@@
for some Markov kernel $\mathsf{K}_{x}^{\beta}$, fully supported on $C_{\beta}(x)$, i.e., $\mathsf{K}_{x}^{\beta}\{C_{\beta}(x)\}=1$ for each $\beta\in[0,1]$ and some probability measure $\mathsf{M}_{x}$ on $[0,1]$ such that a random variable with distribution $\mathsf{M}_{x}$ is stochastically no smaller than ${\sf Unif}(0,1)$.
@@
A relevant question is if there exists a distinguished member $\mathsf{Q}_{x}^{\star}$ of $\mathscr{C}(\overline{\Pi}_{x})$ such that equality on the right-hand side of (7) is achieved for each $\alpha$, i.e.,
@@
There are cases in which attainment of the equality in (9) is out of reach. For instance, suppose that $\theta\mapsto\pi_{x}(\theta)$ is bounded away from 0 for some $x$. This is rare in applications, but not impossible. It may happen that the data is not particularly informative about a feature $\Phi=m(\Theta)$ of $\Theta$, and then the corresponding marginal IM (see Section 4.3) has a non-vanishing contour, i.e., the data is insufficiently informative to exclude any values of $\Phi$ with high degree of confidence. In any case, a probability distribution $\mathsf{Q}_{x}^{\star}$ satisfying (9), if one exists, will be called the inner probabilistic approximation of the possibility measure $\overline{\Pi}_{x}$ (Dubois et al., 2004).
Note that $C_{\alpha}(x)$ in (5) is a $100(1-\alpha)$credible set associated with an inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ and, since the IM’s validity property implies that it’s also a $100(1-\alpha)$confidence set, it follows that $\mathsf{Q}_{x}^{\star}$ exactly—not just asymptotically—achieves the “probability matching” property (e.g., Staicu and Reid, 2008; Datta and Ghosh, 1995) desired of no-prior Bayes solutions.
@@
$\mathsf{Q}_{x}^{\star}$-credible sets are confidence sets.
What does it take to satisfy (9)? From the characterization (8) in Theorem 1, it is relatively easy to see that equality in (9) amounts to a choice of kernel $\mathsf{K}_{x}^{\beta}$ that is fully supported on $\partial C_{\beta}(x)$, the boundary of $C_{\beta}(x)$, and marginal $\mathsf{M}_{x}={\sf Unif}(0,1)$. How the kernel allocates its mass to $\partial C_{\beta}(x)$ isn’t determined by the equality in (9), hence the inner probabilistic approximation of $\overline{\Pi}_{x}$ isn’t unique. But the sets $C_{\beta}(x)$ are typically bounded, i.e., contained in a compact subset of $\mathbb{T}$, and in fact, when the sample size is moderate to large, they’re approximately ellipsoids (Section 4.2). So, in such cases, choosing the kernel $\mathsf{K}_{x}^{\beta}$ to be a uniform distribution on $\partial C_{\beta}(x)$, for each $\beta\in(0,1)$, is quite natural, and consistent with both the maximum entropy (e.g., Cover and Thomas, 2006; Bernardo and Smith, 1994) and indifference principles (e.g., Keynes, 1921; Jaynes, 2003).
To summarize, my proposed inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ of the possibilistic IM is a uniform mixture of uniform distributions. More specifically, a sample $(\Theta\mid x)\sim\mathsf{Q}_{x}^{\star}$ can be drawn via the following two steps:
@@
As a quick illustration, consider a gamma model $\mathsf{P}_{\theta}={\sf Gamma}(n,\theta)$ for a single observation $X$, with $n$ a fixed constant; the more practical case of $n$ iid exponential samples where $X$ is the sum, the complete sufficient statistic, reduces to this one. Then the likelihood-based possibilistic IM has contour function given by
@@
The right-hand side has no simple closed-form expression but, since $X/\theta$ is a pivot under $\mathsf{P}_{\theta}$, the computations are straightforward using Monte Carlo; a complicated exact expression, whose evaluation doesn’t require Monte Carlo, is given in Proposition 1 of Hector et al., (2025). The contour shown in Figure 1 is actually the one above, with $n=7$ and $x=14$. Clearly, the level sets $C_{\alpha}(x)$ are intervals, say, $[a(x,\alpha),b(x,\alpha)]$, where $b\geq a$, but with $a(x,1)=b(x,1)=x/n$, the maximum likelihood estimator. Then the kernel $\mathsf{K}_{x}^{\alpha}$ in the inner probabilistic approximation is a probability distribution supported on the endpoints $a(x,\alpha)$ and $b(x,\alpha)$, i.e.,
@@
where $w_{x,\alpha}\in[0,1]$ and $\delta_{z}$ denotes a probability measure that puts all its mass at $z$. The question I want to address here is: what kind of distribution is an inner probabilistic approximation, what does it look like? In the case where $w_{x,\alpha}\equiv w$ is a fixed value that doesn’t depend on $\alpha$, the distribution function of $\Theta\sim\mathsf{Q}_{x}^{\star}$ can easily be derived:
$\mathsf{Q}_{x}^{\star}(\Theta\leq t)=\begin{cases}w\,\pi_{x}(t)&\text{if $t\leq x/n$}\\ 1-(1-w)\,\pi_{x}(t)&\text{if $t>x/n$}.\end{cases}$
Two such distribution functions are shown in Figure 2(a), one based on $w=0.5$ (solid) and another based on $w\approx 0.45$ (dashed). A third distribution function (dashed) is also shown, namely, the Bayesian posterior distribution function based on a default (right Haar) prior with density $\theta\mapsto\theta^{-1}$, which is an inverse gamma distribution. This Bayes posterior doesn’t obviously have a form like that above, [^2] I wasn’t able to find a constant $w$ that would recover the inverse gamma distribution—the choice $w\approx 0.45$ was my best effort. Consequently, this particular instantiation of the general form of $\mathsf{Q}_{x}^{\star}$ must involve an $\alpha$-dependent weight and it’d be interesting to determine what that $w_{x,\alpha}$ is. but that it too is an inner probabilistic approximation of the IM is a general fact established in Section 4.1 below. The three distribution functions are similar, but not exactly the same. To confirm that all three are, indeed, inner probabilistic approximations, I plot, in Figure 2(b), the non-coverage probability functions $\alpha\mapsto 1-\mathsf{Q}_{x}^{\star}\{C_{\alpha}(x)\}$. According to (9), the non-coverage probability for each inner probabilistic approximation should be equal to $\alpha$, which is confirmed in the plot.
@@
Figure 2: Plots related to three inner probabilistic approximations for the gamma illustration with $n=7$ and $x=14$: uniform kernel (solid), non-uniform kernel with $w\approx 0.45$ (dashed), and the Bayesian posterior distribution (dotted) based on the default prior $\theta\mapsto\theta^{-1}$. Panel (a) shows the cumulative distribution function $t\mapsto\mathsf{Q}_{x}^{\star}(\Theta\leq t)$ and Panel (b) shows the non-coverage probability function $\alpha\mapsto 1-\mathsf{Q}_{x}^{\star}\{C_{\alpha}(x)\}$.

### v3

@@ 3 Inner probabilistic approximations
3.1 Intuition
Section 2.1 discussed approximating a given probability distribution by a possibility measure, the workhorse being the probability-to-possibility transform. The goal of the present section is to reverse this process: to approximate the possibilistic IM by a suitable probability distribution. Before giving a general description, I’ll offer some intuition using a simple example where all the calculations can be done explicitly.
Let $X\sim\mathsf{P}_{\Theta}={\sf N}(\Theta,1)$. Given $X=x$, the standard no-prior Bayes solution, corresponding to the flat, Jeffreys prior, returns $\mathsf{Q}_{x}={\sf N}(x,1)$ as the posterior distribution for $\Theta$. My proposal of approximating the possibilistic IM by a probability definitely isn’t effective if it doesn’t yield the standard no-prior Bayes solution in this simple case. For the IM construction, the relative likelihood is $R(x,\theta)=\exp\{-(x-\theta)^{2}/2\}$ and the contour function is
$\pi_{x}(\theta)$ $=\mathsf{P}_{\theta}\{R(X,\theta)\leq R(x,\theta)\}$
$=\mathsf{P}_{\theta}\{e^{-(X-\theta)^{2}/2}\leq e^{-(x-\theta)^{2}/2}\}$
$=\mathsf{P}\{{\sf ChiSq}(1)\geq(x-\theta)^{2}\}$
$=1-F_{1}\bigl((x-\theta)^{2}\bigr),$
where $F_{\nu}$ is the ${\sf ChiSq}(\nu)$ distribution function. How are $\mathsf{Q}_{x}$ and the corresponding $\mathsf{\Pi}_{x}$ related? One way to see it, following Section 2.1, is to apply the probability-to-possibility transform to $\mathsf{Q}_{x}$ and see what comes out. In this case, with $q_{x}$ the ${\sf N}(x,1)$ density function
$\text{prob-to-poss}_{\mathsf{Q}_{x}}(\theta)$ $:=\mathsf{Q}_{x}\{q_{x}(\Theta)\leq q_{x}(\theta)\}$
$=\mathsf{Q}_{x}\{e^{-(\Theta-x)^{2}/2}\leq e^{-(\theta-x)^{2}/2}\}$
$=\mathsf{P}\{{\sf ChiSq}(1)\geq(\theta-x)^{2}\}$
$=\pi_{x}(\theta).$
That is, the outer possibilistic approximation of $\mathsf{Q}_{x}$, obtained via the probability-to-possibility transform, is precisely the possibilistic IM $\mathsf{\Pi}_{x}$. This is an outer approximation because, roughly, it’s the “least imprecise” possibility measure among those that $\preceq$-dominate $\mathsf{Q}_{x}$. It’s only natural, then, to turn this relationship around and say that $\mathsf{Q}_{x}$ is an inner probabilistic approximation of $\mathsf{\Pi}_{x}$ in the sense of being the “maximally diffuse” probability distribution that’s $\preceq$-dominated by $\mathsf{\Pi}_{x}$. These details will be made more precise in Section 3.2 below.
As a preview, there are at least two equivalent ways to extract a probability from a possibility: one based on bounding tail probabilities and another based on level set matching. Here I’ll focus on the latter approach, formalized in Section 3.2. Since $\pi_{x}(\theta)$ is most naturally interpreted as the “plausibility” of $\Theta=\theta$ given $x$, and since posterior densities are (informally) interpreted in the same way, it makes sense that the density $q_{x}$ of $\mathsf{Q}_{x}$ would have the same level sets as $\pi_{x}$. Akin to slice sampling (e.g., Neal, 2003), a probability $\mathsf{Q}_{x}$ is determined by:
  • first sampling the level set;
  • then sampling a point on the chosen level set.
In the present case, the level sets are the confidence intervals $C_{\alpha}(x)=\{\theta:\pi_{x}(\theta)>\alpha\}$, indexed by $\alpha\in[0,1]$. Then sampling the level sets is equivalent to sampling the level itself, say $A$, from a distribution supported on $[0,1]$. If $\mathsf{Q}_{x}$ represents the distribution of $\Theta$ under the two-stage sampling scheme above, then $\Theta\in C_{\alpha}(x)$ if and only if $A>\alpha$. Then dominance $\mathsf{Q}_{x}\preceq\mathsf{\Pi}_{x}$ implies
$\mathsf{Q}_{x}(A>\alpha)$ $\equiv\mathsf{Q}_{x}\{C_{\alpha}(x)\}$
$=1-\mathsf{Q}_{x}\{C_{\alpha}(x)^{c}\}$
$\geq 1-\mathsf{\Pi}_{x}\{C_{\alpha}(x)^{c}\}$
$=1-\alpha.$
That is, dominance itself $A$ is stochastically no smaller than ${\sf Unif}(0,1)$. At least intuitively, the “maximally diffuse” probability distribution $\mathsf{Q}_{x}^{\star}$ dominated by $\mathsf{\Pi}_{x}$ corresponds to taking $A\sim{\sf Unif}(0,1)$ and, given the level $A=\alpha$, sampling $\Theta$ on $C_{\alpha}(x)$ with maximal variance, i.e., uniformly on the endpoints of $C_{\alpha}(x)$. Those endpoints are $x\pm z_{\alpha/2}$, where $z_{\beta}$ is the upper-$\beta$ quantile of ${\sf N}(0,1)$, and then the distribution function of $\mathsf{Q}_{x}^{\star}$ is
[…]
As a function of $A\sim{\sf Unif}(0,1)$, the random variable $z_{A/2}$ has the same distribution as $|Z|$ with $Z\sim{\sf N}(0,1)$. So,
$\mathsf{Q}_{x}^{\star}\bigl((-\infty,\theta]\bigr)=\tfrac{1}{2}\,\mathsf{P}(|Z|\geq x-\theta)+\tfrac{1}{2}\,\mathsf{P}(|Z|\leq\theta-x).$
Using some basic properties of the standard normal distribution, the right-hand side simplifies, leading to
$\mathsf{Q}_{x}^{\star}\bigl((-\infty,\theta]\bigr)=\mathsf{P}(Z\leq\theta-x),$
and, therefore, this “maximally diffuse” probability distribution $\mathsf{Q}_{x}^{\star}$ dominated by $\mathsf{\Pi}_{x}$ is just the standard no-prior Bayes solution, i.e., $\mathsf{Q}_{x}^{\star}={\sf N}(x,1)$.
3.2 Characterization
@@
The goal of the present section is to formalize and generalize the intuition developed above. Start with a well-known characterization of the credal set for general possibility measures (e.g., Couso et al., 2001; Destercke and Dubois, 2014) which, applied to the present case, says
$\begin{split}\mathsf{Q}_{x}&\in\mathscr{C}(\mathsf{\Pi}_{x})\iff\mathsf{Q}_{x}\{C_{\alpha}(x)\}\geq 1-\alpha\;\;\text{for all $\alpha\in[0,1]$},\end{split}$ (7)
@@
where $C_{\alpha}(x)$ is as defined in (5) with $\pi_{x}$ the contour corresponding to $\mathsf{\Pi}_{x}$. The IM’s validity implies that $C_{\alpha}(x)$ is a $100(1-\alpha)$% confidence set, so, since the elements $\mathsf{Q}_{x}$ of $\mathscr{C}(\mathsf{\Pi}_{x})$ assign probability at least $1-\alpha$ to the $100(1-\alpha)$% confidence set $C_{\alpha}(x)$, there’s good reason to call the elements of $\mathscr{C}(\mathsf{\Pi}_{x})$ confidence distributions. This definition of confidence distributions generalizes those commonly found in the statistics literature (Xie and Singh, 2013; Schweder and Hjort, 2013; Cunen et al., 2020) that—for reasons discussed in Section 4.3—focus mainly on scalar parameter cases.
@@
Refinements to the characterization in (7) can be given. For example, Wasserman, 1990b (, Theorem 2.1) characterizes the credal set associated with a belief function as a collection of suitable mixture distributions. The result below gives a similar mixture characterization for possibility measures, including for the IM output $\mathsf{\Pi}_{x}$.
Theorem 1 (Martin, 2025a ).
@@
Let $\pi_{x}(\theta)$ be the contour associated with the possibility measure $\mathsf{\Pi}_{x}$, and set $C_{\alpha}(x)=\{\theta:\pi_{x}(\theta)>\alpha\}$ as in (5). Then $\mathsf{Q}_{x}\in\mathscr{C}(\mathsf{\Pi}_{x})$ if and only if it can be represented as
$\mathsf{Q}_{x}(\cdot)=\int_{0}^{1}\mathsf{K}_{x}^{\alpha}(\cdot)\,\mathsf{M}_{x}(d\alpha),$ (8)
@@
for some kernel $\mathsf{K}_{x}^{\alpha}$, fully supported on $C_{\alpha}(x)$, so that $\mathsf{K}_{x}^{\alpha}\{C_{\alpha}(x)\}=1$ for each $\alpha\in[0,1]$, and some probability measure $\mathsf{M}_{x}$ on $[0,1]$ whose distribution function is upper-bounded by that of ${\sf Unif}(0,1)$.
The mixture representation (8) of the credal set elements implies that a draw $\Theta$ from $\mathsf{Q}_{x}$ can be obtained via a two-step sampling procedure: draw $A\sim\mathsf{M}_{x}$ and, given $A=\alpha$, draw $\Theta\sim\mathsf{K}_{x}^{\alpha}$. The reader will recognize this two-step procedure from the discussion in Section 3.1. There are, of course, many different choices of $\mathsf{K}_{x}^{\alpha}$ and $\mathsf{M}_{x}$ that can be made, and it’s this ambiguity that makes the credal set a genuine, non-singleton set.
@@
A relevant question is if there exists members $\mathsf{Q}_{x}^{\star}$ of $\mathscr{C}(\mathsf{\Pi}_{x})$ such that equality on the right-hand side of (7) is achieved for each $\alpha$, i.e.,
@@
This formalizes the notion of “maximally diffuse” in Section 3.1. The point is that each element $\mathsf{Q}_{x}$ of the credal set is a confidence distribution in the sense that it assigns $\mathsf{Q}_{x}$-probability at least $1-\alpha$ to $C_{\alpha}(x)$. Then “maximally diffuse” simply means a confidence distribution that’s not conservative, one whose probability exactly matches the confidence level. As explained in Section 3.1 above, I’ll refer to those $\mathsf{Q}_{x}^{\star}$ for which (9) holds as inner probabilistic approximations of the possibilistic IM $\mathsf{\Pi}_{x}$.
Setting aside the question about existence of an inner probabilistic approximation for the moment, note that the mixture formulation in (8) implies that the highest posterior density credible sets exactly agree with $\{C_{\alpha}(x):\alpha\in[0,1]\}$ in (5). Since validity of the IM implies that the latter are exact confidence sets, then the former are too. That is, inner probabilistic approximations exactly—not just asymptotically—achieve the “probability matching” property (e.g., Staicu and Reid, 2008; Datta and Ghosh, 1995) desired of no-prior Bayes solutions.
@@
If $\mathsf{Q}_{x}^{\star}$ is an inner probabilistic approximation of $\mathsf{\Pi}_{x}$, then its associated highest posterior density credible sets are confidence sets.
There are cases in which attainment of the equality in (9) is out of reach. For instance, suppose that $\theta\mapsto\pi_{x}(\theta)$ is bounded away from 0 for some $x$. This is rare in applications, but not impossible. It may happen that the data is not particularly informative about a feature $\Phi=m(\Theta)$ of $\Theta$, and then the corresponding marginal IM (see Section 4.3) has a non-vanishing contour—that is, the data is insufficiently informative to exclude any values of $\Phi$ with high degree of confidence.
Some clues about what it takes to satisfy (9) were already presented in Section 3.1. Indeed, to get a “maximally diffuse” member of the credal set, I suggested taking the marginal distribution $\mathsf{M}_{x}$ to be ${\sf Unif}(0,1)$ and taking the kernel $\mathsf{K}_{x}^{\alpha}$ to be fully supported on the boundary, $\partial C_{\alpha}(x)$, of the level set $C_{\alpha}(x)$. To justify this, note that, for each $\alpha\in[0,1]$, a $\mathsf{Q}_{x}$ of the form (8) satisfies
$\mathsf{Q}_{x}\{C_{\alpha}(x)\}$ $=\int_{0}^{1}\mathsf{K}_{x}^{\beta}\{C_{\alpha}(x)\}\,\mathsf{M}_{x}(d\beta)$
$=\int_{0}^{\alpha}\underbrace{\mathsf{K}_{x}^{\beta}\{C_{\alpha}(x)\}}_{\geq 0}\,\mathsf{M}_{x}(d\beta)+\int_{\alpha}^{1}\underbrace{\mathsf{K}_{x}^{\beta}\{C_{\alpha}(x)\}}_{=1}\,\mathsf{M}_{x}(d\beta)$
$\geq\mathsf{M}_{x}([\alpha,1])$
$\geq 1-\alpha.$
Then it’s clear that equality in (9) is achieved by taking $\mathsf{M}_{x}={\sf Unif}(0,1)$ and $\mathsf{K}_{x}^{\alpha}$ fully supported on $\partial C_{\alpha}(x)$, so that $\mathsf{K}_{x}^{\beta}\{C_{\alpha}(x)\}=0$ for all $\beta<\alpha$.
How the kernel allocates its mass to $\partial C_{\alpha}(x)$ isn’t determined by (9), hence the inner probabilistic approximation of $\mathsf{\Pi}_{x}$ isn’t unique. I’ll say more about the role played by the kernel later and in Appendix B. Here I present my recommended choice of inner probabilistic approximation. Note that the sets $C_{\alpha}(x)$ are typically bounded; in fact, when the sample size is moderate to large, they’re approximately ellipsoids (Section 4.2). Then choosing the kernel $\mathsf{K}_{x}^{\alpha}$ to be a uniform distribution on $\partial C_{\alpha}(x)$, for each $\alpha\in(0,1)$, is quite natural, and consistent with both the maximum entropy (e.g., Cover and Thomas, 2006; Bernardo and Smith, 1994) and indifference principles (e.g., Keynes, 1921; Jaynes, 2003). It’ll also be seen in Section 4.2 that uniformity is needed for asymptotic efficiency of the inner probabilistic approximation.
To summarize, my proposed inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ of $\mathsf{\Pi}_{x}$ is a uniform mixture of uniform distributions. More specifically, a sample $(\Theta\mid x)\sim\mathsf{Q}_{x}^{\star}$ can be drawn via the following two-step procedure:
@@
3.3 Illustration
Consider a gamma model $\mathsf{P}_{\theta}={\sf Gamma}(n,\theta)$ for a single observation $X$, with $n$ a fixed constant; the more practical case of $n$ iid exponential samples where $X$ is the sum, the complete sufficient statistic, reduces to this one. Then the likelihood-based IM has contour
@@ 3.3 Illustration
The right-hand side has no simple closed-form expression but, since $X/\theta$ is a pivot under $\mathsf{P}_{\theta}$, the computations are straightforward using Monte Carlo. The contour shown in Figure 1 is actually the one above, with $n=7$ and $x=14$. Clearly, the level sets $C_{\alpha}(x)$ are intervals, say, $a(x,\alpha),b(x,\alpha)]$, where $b\geq a$, but with $a(x,1)=b(x,1)=x/n$, the maximum likelihood estimator. Then the kernel $\mathsf{K}_{x}^{\alpha}$ associated with an inner probabilistic approximation is a probability distribution supported on the endpoints $a(x,\alpha)$ and $b(x,\alpha)$, i.e.,
@@
where $w_{x,\alpha}\in[0,1]$ and $\delta_{z}$ denotes a probability measure that puts all its mass at $z$. The question I want to address here is: what do the inner probabilistic approximations $\mathsf{Q}_{x}^{\star}$ look like? With constant weights, i.e., $w_{x,\alpha}\equiv w$, the argument used in the Gaussian example above can be generalized as follows. For a generic $\theta$, it follows that
$\{a(x,\alpha)\geq\theta\}\;\text{or}\;\{b(x,\alpha)\leq\theta\}\iff\pi_{x}(\theta)\leq\alpha.$
Using the mixture characterization in (8), the distribution function of $\Theta\sim\mathsf{Q}_{x}^{\star}$, at $\theta$ is
$\mathsf{Q}_{x}^{\star}\bigl((-\infty,\theta]\bigr)$ $=\int_{0}^{1}w\,1\{a(x,\alpha)\leq\theta\}\,d\alpha+\int_{0}^{1}(1-w)\,1\{b(x,\alpha)\leq\theta\}\,d\alpha$
$=\begin{cases}w\,\pi_{x}(\theta)&\text{if $\theta\leq x/n$}\\ 1-(1-w)\,\pi_{x}(\theta)&\text{if $\theta>x/n$}.\end{cases}$
Two such distribution functions are shown in Figure 2(a), one based on my recommended choice $w=0.5$ (solid) and another based on $w\approx 0.45$ (dashed). A third distribution function (dotted) is also shown, namely, that of the Bayesian posterior based on the right Haar prior with density $\theta\mapsto\theta^{-1}$, which is an inverse gamma distribution. Unlike the flat (right Haar) prior Bayes posterior in the Gaussian example above, this Bayes posterior can’t be expressed via a constant-weight kernel (Appendix B); the choice $w\approx 0.45$ was my best effort to do so. But that this Bayes solution is also an inner probabilistic approximation is a general fact established in Section 4.1 below. The three distribution functions are very similar, but not the same. To confirm that all three are, indeed, inner probabilistic approximations, I plot, in Figure 2(b), the non-coverage probability functions $\alpha\mapsto 1-\mathsf{Q}_{x}^{\star}\{C_{\alpha}(x)\}$. According to (9), the non-coverage probability for each inner probabilistic approximation should be equal to $\alpha$, which is confirmed in the plot.
@@
Figure 2: Plots of three inner probabilistic approximations for the gamma illustration with $n=7$ and $x=14$: uniform kernel (solid), non-uniform kernel with $w\approx 0.45$ (dashed), and the Bayesian posterior distribution (dotted) based on the default prior $\theta\mapsto\theta^{-1}$. Panel (a) shows the distribution function $t\mapsto\mathsf{Q}_{x}^{\star}(\Theta\leq t)$ and Panel (b) shows the non-coverage probability function $\alpha\mapsto 1-\mathsf{Q}_{x}^{\star}\{C_{\alpha}(x)\}$.